# Uncovering Treatment Effect Heterogeneity in Pragmatic Gerontology Trials

### v1

@@ 4 Results
Posterior summaries of the stratum proportions and the overall CSACE are reported in Table 3.
@@
Figure 2 displays the distribution of individual CSACE estimates among likely always-survivors. The left panel shows sorted posterior means with 95% CrIs; nearly all intervals cover zero, indicating limited evidence for large individual effects, although the spread of point estimates suggests heterogeneity. The right panel histogram is approximately centered at zero. The overall mean CSACE was 0.57-0.57 (95% CrI [3.58, 2.53][-3.58,\,2.53]), indicating no compelling difference between telecare and usual care in the 12 month EQ-5D–VAS summary score for the average always-survivor.
Figure 2: Distribution of CSACE on EQ-5D–VAS at 12 months among likely always-survivors. Left: sorted posterior mean CSACEs with 95% CrIs. Right: histogram of posterior mean CSACEs.
To explore effect modification among likely always-survivors, we adopted a Bayesian “fit-the-fit” strategy: hahn2020bayesian posterior mean CSACEs were used as responses in a classification and regression tree (CART) fitted to baseline covariates. The resulting tree (Figure 3) uses baseline deprivation score for the primary split at 15. Participants with deprivation $\geq 15$ had a mean CSACE of $0.61$ (95% CrI [2.66, 4.07][-2.66,\,4.07]), whereas those with deprivation <15<15 had a mean CSACE of 4.20-4.20 (95% CrI [9.23, 0.72][-9.23,\,0.72]), suggesting a greater tendency toward outcomes favoring usual care in the latter group.
@@
Among participants with low deprivation (<15<15), the tree highlighted the joint role of baseline mental health and baseline EQ–5D–VAS. In particular, because higher scores indicate better baseline status, those with poor baseline mental health (score <32<32) and low baseline quality of life (EQ–5D–VAS <28<28) showed a mean CSACE of 9.00-9.00 (95% CrI [19.77, 0.58][-19.77,\,0.58]), favoring usual care in that subgroup. Overall, the tree points to higher deprivation—especially in non–alone households—as associated with greater benefit from telecare, whereas low deprivation combined with poorer mental health and lower baseline EQ-5D–VAS is associated with harm. Because many terminal node intervals include zero, these subgroup findings should be interpreted as exploratory.
Figure 3: CART applied to posterior mean CSACEs for EQ-5D–VAS at 12 months among likely always-survivors. Nodes display posterior mean CSACE and 95% CrI for each subgroup; splits are on baseline covariates.
@@ 5 Discussion
This study develops and demonstrates the application of machine learning tools in identifying treatment effect heterogeneity in a pragmatic gerontological trial. Specifically, we re-examine the WSD Telecare Questionnaire Study using a principal stratification causal framework for cluster-randomized trials with outcomes truncated by death. We embed mixed-effects BART to model strata membership and arm and stratum specific outcomes, yielding estimates of the SACE and individual survival treatment effect while accounting for cluster dependence and flexible nonlinear associations between baseline covariates and outcomes. Three findings are worth highlighting.
@@
First, the mean CSACE on EQ–5D–VAS at 12 months was close to zero, indicating no compelling average improvement among always–survivors. Because SACE conditions on principal strata rather than observed survival, this estimand differs conceptually from intent-to-treat and survivor only analyses. Our result suggests that, for individuals who would survive regardless of arm, telecare did not yield apparent improvements in overall quality of life as captured by EQ–5D–VAS.
@@
Second, the heterogeneity analysis among likely always–survivors points to clinically meaningful variation. The most important baseline features that appear to moderate the response to telecare include community deprivation, household structure, and mental health (MS-12). Deprivation plausibly indexes environmental and social risks as well as unmet safety needs; when deprivation is higher, passive monitoring and escalation pathways have greater scope to prevent adverse events or provide reassurance, whereas in more advantaged settings the incremental value of monitoring may be limited and, for some, burdensome. Living arrangement modifies this gradient: in households with co-residents, reminders, help with equipment, interpretation of alerts, and faster response can translate monitoring into tangible benefit, while individuals living alone may not realize the same gains. Finally, poorer baseline mental health and low baseline EQ–5D–VAS mark greater symptom burden and psychosocial stress; passive monitoring alone may not address these drivers of low self-rated health and may heighten vigilance without relief, aligning with the negative signals observed in those branches. Taken together, these patterns argue for targeted deployment—prioritizing high deprivation recipients with household support while treating the subgroup signals as exploratory and requiring confirmation in pre-specified analyses. Meanwhile, it is worth highlighting that credible intervals at most terminal nodes include zero and subgroup sizes are modest; these findings should be regarded as hypothesis generating and framed as guidance for future, pre-specified confirmatory analyses rather than as definitive subgroup effects.
@@
Third, the proposed estimation strategy handled multiple sources of incompleteness common in gerontology trials. The nested–MAR assumption separates missing survival status from, conditional on observed survival, missing outcomes, thereby permitting ignorable likelihood–based inference while retaining a clear estimand. Coupled with mixed–effects BART, the approach balances design based features of CRTs (cluster randomization, within–cluster dependence) with nonparametric flexibility in the strata and outcome models, reducing vulnerability to functional form misspecification relative to linear mixed models.
@@
Our results relies on a series of structural assumptions. At the trial design level, we adopt SUTVA with partial interference, meaning outcomes depend only on a person’s cluster assignment, with cluster-level randomization and positivity ensuring independence of assignment and potential outcomes; in the WSD setting, where randomization occurred at the practice level and telecare was delivered within households, these conditions are reasonable, although some unobserved cross-practice interactions cannot be ruled out. At the modeling layer, we use a principal stratification framework with monotonicity for survival ($S_{ij}(1)\geq S_{ij}(0)$), which reflects the intervention’s intent, accommodate the outcomes are truncated by death, and rely on principal exchangeability, an assumption that conditional on observed covariates, models for strata membership and outcomes are sufficiently rich to capture the relevant variation. While these conditions cannot be verified directly, they are standard in mixture-based analyses and provide a defensible path to identification. For missingness, we assume a nested MAR structure, where survival status and outcomes are missing at random given treatment and covariates, which is plausible given the WSD data and reasons for follow-up loss, though we note developing methods to relax this assumption would further strengthen robustness. Finally, we recognize that the analysis of effect modification is a post hoc exploration of potential mechanism rather than a confirmatory. Nevertheless, the results highlight directions for future work and potential target population and improvement in the intervention design. Overall, our study shows that principal stratification combined with mixed effects BART yields inference that respects the trial design and is robust to nonlinear relations and higher order interactions. Substantively, the WSD reanalysis suggests that broad telecare deployment may not improve average quality of life among always–survivors; focusing implementation on participants in higher deprivation settings and those living with others may be more beneficial.
@@ 6 Conclusion
Two messages stand out. First, on average, telecare did not improve the 12 month EQ–5D–VAS among people who would survive regardless of assignment. Second, there are meaningful differences across subgroups. Signals of benefit were more apparent for participants living in more deprived settings, especially those residing with others who could support consistent engagement with the system. In contrast, participants with lower deprivation combined with poorer baseline mental health and low baseline EQ–5D–VAS showed little evidence of benefit and may do better with usual care. These patterns are exploratory, but they suggest a practical strategy: target telecare to patients and households with greater capacity to translate monitoring into action.
For clinicians and service planners, the implication is to move away from uniform deployment strategies. Prioritizing more deprived patients and ensuring household or caregiver support may yield greater gains, whereas those with substantial psychological burden and low baseline quality of life may need complementary services (e.g., mental health support, rehabilitation) alongside any monitoring technology. For researchers, future trials should prospectively evaluate such targeting rules, collect baseline factors that plausibly modify benefit, and pre-specify confirmatory subgroup analyses. The same analytical framework applies directly to other health-related quality of life outcomes collected in this study, including the SF-12 physical and mental component scores, enabling a broader view of patient-centered benefit. Replication in other settings, sensitivity analyses to key assumptions, and linkage to cost effectiveness will further clarify when and for whom telecare improves quality of life in later life.
@@ Acknowledgments
Research in this article was partially supported by the Patient-Centered Outcomes Research Institute® (PCORI® Award ME-2020C1-19220), the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (grant numbers R01-HL168202, R01-HL178513), as well as by the National Institute on Aging (NIA) of the National Institutes of Health (NIH) under Award Number U54AG063546-06A1, which funds NIA Imbedded Pragmatic Alzheimer’s Disease (AD) and AD-Related Dementias Clinical Trials Collaboratory (NIA IMPACT Collaboratory), the Yale OAIC (P30AG021342), and the Yale ADRC (P30AG066508). All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH or PCORI® or its Board of Governors or Methodology Committee. The Whole Systems ‘Demonstrator’ (WSD) telecare trial study was funded by the Policy Research Programme in the Department of Health, UK. The views expressed are not necessarily those of the Department. The WSD study was approved by Liverpool Research Ethics Committee (ref: 08/H1005/4). We acknowledge the generous sharing of the de-identified WSD telecare trial data from Drs Shashivadan P Hirani and Stanton P Newman at City, University of London, UK.

### v2

@@ 5 Demonstration of the WSD Telecare Questionnaire Study
Posterior summaries of the stratum proportions and the overall SACE are reported in Table 5.
@@
Figure 2 summarizes the distribution of individual CSACE estimates among likely always-survivors, defined as participants with posterior probability at least 0.80 of belonging to the always-survivor principal stratum. In the left panel, each position on the horizontal axis corresponds to a single participant, and the solid curve shows the posterior mean CSACE values sorted from smallest to largest. The light band around the curve represents 95% credible intervals for each individual effect. Across patients, the posterior means range from modest negative to modest positive values, indicating meaningful heterogeneity in the estimated effects, but nearly all intervals include zero, suggesting limited evidence for large individual-level benefits or harms. The right panel displays a histogram of the posterior mean CSACEs with a smooth density overlay. The distribution is approximately symmetric and centered slightly below zero. The overall mean CSACE was 0.57-0.57 (95% CrI [3.58, 2.53][-3.58,\,2.53]), indicating no compelling difference between telecare and usual care in the 12 month EQ-5D-VAS summary score for the average always-survivor.
Figure 2: Distribution of CSACE on EQ-5D-VAS at 12 months among likely always-survivors. Left: participant-level posterior mean CSACEs (sorted) with 95% credible intervals. Right: histogram of posterior mean CSACEs with a smoothed density curve.
To explore effect modification among likely always-survivors, we adopted a Bayesian “fit-the-fit” strategy: 15 posterior mean CSACEs were used as responses in a classification and regression tree (CART) fitted to baseline covariates. The resulting tree (Figure 3) uses baseline deprivation score for the primary split at 15. Participants with deprivation $\geq 15$ had a mean CSACE of $0.61$ (95% CrI [2.66, 4.07][-2.66,\,4.07]), whereas those with deprivation <15<15 had a mean CSACE of 4.20-4.20 (95% CrI [9.23, 0.72][-9.23,\,0.72]), suggesting a greater tendency toward outcomes favoring usual care in the latter group.
@@
Among participants with low deprivation (<15<15), the tree highlighted the joint role of baseline mental health and baseline EQ-5D-VAS. In particular, because higher scores indicate better baseline status, those with poor baseline mental health (score <32<32) and low baseline quality of life (EQ-5D-VAS <28<28) showed a mean CSACE of 9.00-9.00 (95% CrI [19.77, 0.58][-19.77,\,0.58]), favoring usual care in that subgroup. Overall, the tree points to higher deprivation—especially in non–alone households—as associated with greater benefit from telecare, whereas low deprivation combined with poorer mental health and lower baseline EQ-5D-VAS is associated with harm. Because many terminal node intervals include zero, these subgroup findings should be interpreted as exploratory.
Figure 3: CART applied to posterior mean CSACEs for EQ-5D-VAS at 12 months among likely always-survivors. Nodes display posterior mean CSACE and 95% CrI for each subgroup; splits are on baseline covariates.
@@ 6 Discussion
This study develops and demonstrates the application of machine learning tools in identifying treatment effect heterogeneity in a pragmatic gerontological trial. Specifically, we re-examine the WSD Telecare Questionnaire Study using a principal stratification causal framework for cluster-randomized trials with outcomes truncated by death. We embed mixed-effects BART to model strata membership and arm- and stratum-specific outcomes, yielding estimates of the SACE and individual survival treatment effect while accounting for cluster dependence and flexible nonlinear associations between baseline covariates and outcomes. Three findings are worth highlighting.
@@
First, the mean CSACE on EQ-5D-VAS at 12 months was close to zero, indicating no compelling average improvement among always-survivors. Because SACE conditions on principal strata rather than observed survival, this estimand differs conceptually from intent-to-treat and survivor only analyses. Our result suggests that, for individuals who would survive regardless of arm, telecare did not yield apparent improvements in overall quality of life as captured by EQ-5D-VAS.
@@
Second, the heterogeneity analysis among likely always-survivors points to clinically meaningful variation. The most important baseline features that appear to moderate the response to telecare include community deprivation, household structure, and mental health. Deprivation plausibly indexes environmental and social risks as well as unmet safety needs; when deprivation is higher, passive monitoring and escalation pathways have greater scope to prevent adverse events or provide reassurance, whereas in more advantaged settings the incremental value of monitoring may be limited and, for some, burdensome. Living arrangement modifies this gradient: in households with co-residents, reminders, help with equipment, interpretation of alerts, and faster response can translate monitoring into tangible benefit, while individuals living alone may not realize the same gains. Finally, poorer baseline mental health and low baseline EQ-5D-VAS mark greater symptom burden and psychosocial stress; passive monitoring alone may not address these drivers of low self-rated health and may heighten vigilance without relief, aligning with the negative signals observed in those branches. Taken together, these patterns argue for targeted deployment—prioritizing high deprivation recipients with household support while treating the subgroup signals as exploratory and requiring confirmation in pre-specified analyses. Meanwhile, it is worth highlighting that credible intervals at most terminal nodes include zero and subgroup sizes are modest; these findings should be regarded as hypothesis generating and framed as guidance for future, pre-specified confirmatory analyses rather than as definitive subgroup effects.
@@
Third, the proposed estimation strategy handled multiple sources of incompleteness common in gerontology trials. The nested MAR assumption separates missing survival status from, conditional on observed survival, missing outcomes, thereby permitting ignorable likelihood–based inference while retaining a clear estimand. Coupled with mixed–effects BART, the approach balances design based features of CRTs (cluster randomization, within–cluster dependence) with nonparametric flexibility in the strata and outcome models, reducing vulnerability to functional form misspecification relative to linear mixed models.
@@
Our results rely on a series of structural assumptions. At the trial design level, we adopt SUTVA with partial interference, meaning outcomes depend only on a person’s cluster assignment, with cluster-level randomization and positivity ensuring independence of assignment and potential outcomes; in the WSD setting, where randomization occurred at the practice level and telecare was delivered within households, these conditions are reasonable, although some unobserved cross-practice interactions cannot be ruled out. At the modeling layer, we use a principal stratification framework with monotonicity for survival ($S_{ij}(1)\geq S_{ij}(0)$), which reflects the intervention’s intent, accommodate outcomes truncated by death, and rely on principal exchangeability, an assumption that conditional on observed covariates, models for strata membership and outcomes are sufficiently rich to capture the relevant variation. While these conditions cannot be verified directly, they are standard in mixture-based analyses and provide a defensible path to identification. For missingness, we assume a nested MAR structure, where survival status and outcomes are missing at random given treatment and covariates, which is plausible given the WSD data and reasons for follow-up loss, though we note developing methods to relax this assumption would further strengthen robustness. Finally, we recognize that the analysis of effect modification is a post hoc exploration of potential mechanism rather than confirmatory. Nevertheless, the results highlight directions for future work and potential target population and improvement in the intervention design. Overall, our study shows that principal stratification combined with mixed effects BART yields inference that respects the trial design and is robust to nonlinear relations and higher order interactions. Substantively, the WSD reanalysis suggests that broad telecare deployment may not improve average quality of life among always-survivors; focusing implementation on participants in higher deprivation settings and those living with others may be more beneficial.
@@ 7 Conclusion
Two messages stand out. First, on average, telecare did not improve the 12 month EQ-5D-VAS among people who would survive regardless of assignment. Second, there are meaningful differences across subgroups. Signals of benefit were more apparent for participants living in more deprived settings, especially those residing with others who could support consistent engagement with the system. In contrast, participants with lower deprivation combined with poorer baseline mental health and low baseline EQ-5D-VAS showed little evidence of benefit and may do better with usual care. These patterns are exploratory, but they suggest a practical strategy: target telecare to patients and households with greater capacity to translate monitoring into action.
From a translational perspective, several practical lessons emerge. Our analysis relies on a monotonicity assumption that telecare does not worsen survival compared with usual care. In many late-phase trials, where interventions have undergone safety piloting and are not expected to increase mortality, this is a reasonable working assumption; when there is genuine concern about treatment-induced harm, the framework can in principle be extended to include a “harmed” principal stratum, at the cost of substantially more complex mixture modeling and larger sample-size requirements to identify a typically rare subgroup. Likewise, the nested MAR assumption for missing survival status and quality-of-life outcomes is most credible when investigators collect rich baseline prognostic measures and understand the reasons for missingness in detail. In applications, we recommend that analysts (i) carefully document patterns and causes of missing data, (ii) assess the plausibility of nested MAR given the available covariates, and (iii) use simulation-based sensitivity analyses to gauge the impact of deviations from these assumptions. Finally, mixed-effects BART is particularly advantageous when there is a non-negligible intracluster correlation and a moderate to large number of clusters, whereas in settings with very small ICC or few clusters, simpler cluster-agnostic BART models may perform competitively and can be considered as a pragmatic alternative.
For clinicians and service planners, the implication is to move away from uniform deployment strategies. Prioritizing more deprived patients and ensuring household or caregiver support may yield greater gains, whereas those with substantial psychological burden and low baseline quality of life may need complementary services (e.g., mental health support, rehabilitation) alongside any monitoring technology. For researchers, future trials should prospectively evaluate such targeting rules, collect baseline factors that plausibly modify benefit, and pre-specify confirmatory subgroup analyses. The same analytical framework can be applied to other health-related quality of life outcomes collected in this study, including the SF-12 physical and mental component scores, enabling a broader view of patient-centered benefit. Replication in other settings, sensitivity analyses to key assumptions, and linkage to cost effectiveness will further clarify when and for whom telecare improves quality of life in later life.
@@ Acknowledgments
Research in this article was partially supported by the Patient-Centered Outcomes Research Institute® (PCORI® Award ME-2020C1-19220), the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (grant numbers R01-HL168202, R01-HL178513), as well as by the National Institute on Aging (NIA) of the National Institutes of Health (NIH) under Award Number U54AG063546-06A1, which funds NIA Imbedded Pragmatic Alzheimer’s Disease (AD) and AD-Related Dementias Clinical Trials Collaboratory (NIA IMPACT Collaboratory), the Yale OAIC (P30AG021342), and the Yale ADRC (P30AG066508). All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH or PCORI® or its Board of Governors or Methodology Committee. The Whole Systems ‘Demonstrator’ (WSD) telecare questionnaire study was funded by the Policy Research Programme in the Department of Health, UK. The views expressed are not necessarily those of the Department. The WSD study was approved by Liverpool Research Ethics Committee (ref: 08/H1005/4). We acknowledge the generous sharing of the de-identified WSD telecare trial data from Drs Shashivadan P Hirani and Stanton P Newman at City, University of London, UK.